You have quarters and dimes that total $2.80. Your friend says it is possible that the number of quarters is 8 more than the number of dimes. Is your friend correct? Explain. Let d represent the number of dimes.
An equation that models this situation is ------------------=2.80.

Answer by jorel1380(3719)

Let d be the number of dimes. Then the number of quarters would be d+8. So:
10d+25(d+8)=280
35d+200=280
35d=80
d=80/35
Since you can't have a fraction of a dime, your friend is incorrect.
☺☺☺☺